Let Timco use a capital structure that is 30% debt and 70% equity, The firm can borrow at 6%. The tax rate is 40%. Timco is planning on paying dividend of 2 to maintain our 5% growth. The current stock price is 16.13. Find the WACC.
Answer #1

After-tax cost of debt=6*(1-tax rate)

=6(1-0.4)=3.6%

Cost of equity=(D1/Current price)+Growth rate

=(2/16.13)+0.05

=17.399256%(Approx).

WACC=Respective costs*Respective weight

=(3.6*0.3)+(0.7*17.399256)

which is equal to

=13.26%(Approx).